# Text2SQL-Flow: A Robust SQL-Aware Data Augmentation Framework for Text-to-SQL

### v2

@@ Abstract
The data-centric paradigm has emerged as a pivotal direction in artificial intelligence (AI), relying on high-quality training data. This shift is especially critical in the Text-to-SQL task, where model performance is constrained by the scarcity, limited diversity, and structural simplicity of existing datasets. To address these challenges, we propose Text2SQL-Flow, a SQL-aware data augmentation framework that systematically generates large-scale, semantically valid, and structurally diverse Text-to-SQL pairs from limited seed data. Our framework operates along six augmentation dimensions and integrates an end-to-end pipeline featuring SQL execution verification, natural language (NL) question generation, chain-of-thought (CoT) reasoning trace generation, and data classification. A modular Database Manager further ensures cross-database compatibility and scalability. Leveraging this framework, we construct SQLFlow, a high-quality dataset comprising 89,544 annotated examples. We demonstrate the utility of SQLFlow in both fine-tuning and prompt-based settings: (1) For open-source large language models (LLMs), fine-tuning with SQLFlow enhances the problem-solving capabilities. Under the same data budget, models trained on SQLFlow achieve competitive performance gains across multiple benchmarks. (2) For closed-source LLMs, we propose a masked alignment retrieval method that leverages SQLFlow as both a knowledge base and the training data for the retrieval model. This approach enables structure-aware example matching by modeling fine-grained alignments between NL questions and SQL queries. Experimental results show that our retrieval strategy outperforms existing example retrieval methods, highlighting the dual importance of SQLFlow’s high-quality data and our novel retrieval technique. Our work establishes a scalable, data-centric foundation for advancing Text-to-SQL systems and underscores the indispensable role of structured, high-fidelity data in modern AI development. Our code is available at https://github.com/TechNomad-ds/Text2SQL-Flow.
@@ I Introduction
In recent years, the data-centric artificial intelligence (AI) paradigm has garnered increasing attention [zha2025data, jakubik2024data]. Traditional algorithm-centric approaches primarily focus on expanding model architectures and optimizing learning algorithms. However, in many cutting-edge fields, the main bottleneck of development has gradually shifted from algorithmic complexity to the availability of high-quality data. Continuous optimization of algorithms is facing diminishing marginal returns, while vast amounts of data remain underutilized, containing immense potential value. Therefore, the data-centric perspective emphasizes the core role of high-quality, diverse, and well-structured training data in advancing artificial intelligence [haukkala2022data].
Taking large language models (LLMs) as an example, their generalization ability and robustness highly depend on the breadth and quality of the training data. Similarly, in downstream tasks such as domain adaptation, high-quality data can serve both as reference material for generating answers and as guidance for solving problems [wang2024survey]. Consequently, building scalable and high-quality datasets has become a cornerstone for advancing artificial intelligence.
Figure 1: Performance comparison of models trained on SQLFlow vs. SynSQL [li2025omnisql] (both at 89,544 samples). Blue labels indicate results for fine-tuned open-source models, while green labels indicate results of closed-source model with few-shot retrieval for prompt construction.
@@
Window Func. Set Op. Subquery Aggregation # CASE # WHERE # Join
Spider-train 8,659 0.00 6.07 15.59 22.6 0.00 0.86 0.71
SQLFlow-Spider 37,517 5.27 7.37 25.55 40.75 0.19 1.50 1.04
@@
SQLFlow-EHRSQL 14,491 9.22 4.11 48.13 33.43 0.17 3.56 1.03
SQLFlow 89,544 5.37 5.77 27.19 34.80 0.20 1.97 1.09
@@
The Text-to-SQL task is currently facing an urgent demand for high-quality data. Although the structured query language (SQL) enables powerful and precise database interactions, its complex syntax and steep learning curve limit its accessibility for non-expert users. The Text-to-SQL task aims to automatically translate natural language (NL) questions into executable SQL queries, allowing users without specialized knowledge to access databases conveniently [kanburouglu2024text]. This capability shows broad application prospects in intelligent analysis, business intelligence platforms, and enterprise data services. With the rise of LLMs, significant progress has been made in Text-to-SQL research. Existing methods are mainly divided into two categories: (1) fine-tuning-based methods [gao2023text, gao2024xiyan, li2024codes, sheng2025slm] and (2) prompt-based methods [wei2022chain, tai2023exploring, li2023resdsql, pourreza2024din]. Although these methods have competitive performance on mainstream benchmarks, their further development is still constrained by the scarcity of high-quality data.
@@
In both categories of methods, high-quality data plays a pivotal role within the data-centric AI paradigm. Fine-tuning-based methods rely on data pairs to enable LLMs to learn the mapping from NL questions to SQL queries, thereby enhancing the model’s reasoning capability. In prompt-based methods, where model weights are inaccessible, in-context learning becomes key to improving performance, and the quality of retrieved few-shot examples is critical. Data can serve as reference examples in a knowledge base or be used to train retrieval models for obtaining more representative few-shot samples [liu2023comprehensive, nan2023enhancing, gao2023text].
Since the Text-to-SQL task primarily relies on data pairs composed of NL questions and corresponding SQL queries, efficiently obtaining these data pairs becomes a critical challenge. Manual collection not only struggles to guarantee quality but is also limited in scale. Existing public datasets (e.g., Spider [yu2018spider], BIRD [li2024can]) mainly rely on manual annotation, suffering from limited size, poor scalability, and narrow data coverage. Recent studies (e.g., OmniSQL [li2025omnisql]) have attempted to explore synthetic data generation pathways. However, such methods are often costly, generate SQL with excessive variation, and require large data volumes to achieve performance improvements. Thus, achieving scalable, high-quality data efficiently has become a central challenge.
To address the aforementioned issues, we propose Text2SQL-Flow, a robust Text-to-SQL data augmentation framework. This framework can efficiently augment original Text-to-SQL data, enhance data diversity, and possesses good scalability. Unlike synthesizing data from scratch, our framework builds upon existing SQL queries, defining six data augmentation dimensions covering Data Value Transformations, Query Structure Modifications, and Complexity Enhancements, among others, utilizing LLMs to synthesize diverse and semantically valid augmented SQL data. A data synthesis pipeline is designed to automate the generation of high-quality data. Various operators are responsible for SQL execution filtering, corresponding NL question generation, prompt generation, chain-of-thought (CoT) path generation for problem-solving, and data classification operators that categorize the generated augmented data to construct high-quality datasets. This framework can automatically generate large-scale SQL variants with diverse stylistic patterns, building high-quality, large-scale datasets from originally small-scale data. Furthermore, to improve the framework’s adaptability to different databases and enhance database operation efficiency, a Database Manager module is designed. It interacts with various databases through a Database Connector interface. Each database inherits from and implements the methods defined in the Database Connector, while the Database Manager performs high-level invocations, thereby enhancing extensibility across different databases.
@@
Based on our framework, a high-quality Text-to-SQL dataset comprising 89,544 entries called SQLFlow is constructed. Using SQLFlow, we explore the applications of enhanced data in Text-to-SQL tasks. For open-source LLMs, the augmentation framework can provide higher-quality training data. The included CoT can further assist models in improving their reasoning capabilities and problem-solving proficiency. After fine-tuning open-source models with data generated by our framework, we achieved excellent performance on multiple benchmarks, validating the framework’s effectiveness. As shown in Figure 1, under the same data scale, the model trained by SQLFlow achieves higher performance than SynSQL [li2025omnisql] across all benchmarks, attaining 59.2% on the BIRD-dev dataset and 83.5% on the Spider-test dataset. For closed-source models, we adopt prompt-based methods, retrieving few-shot examples to provide references for the LLM in solving problems, thereby achieving broad applicability across various scenarios. Existing systems typically select examples from limited public datasets, which have significant limitations in semantic diversity and structural coverage. Therefore, the example knowledge base for retrieval can also be constructed using the data generated by our framework. However, at the retrieval method level, current mainstream retrieval strategies often focus merely on surface-level semantic similarity, neglecting the deep syntactic and logical relationships between NL questions and SQL queries. To address this, we propose an improved masked alignment retrieval method for example selection. This method learns the masked alignment relationships between NL questions and SQL queries during the training phase and achieves fine-grained structure-aware matching during the inference phase, enhancing the relevance and effectiveness of retrieved examples. The training data for this retrieval model also utilizes the data constructed by our framework. As shown in Figure 1 shows that our retrieval method achieves competitive performance on benchmarks, with 61.0% on the BIRD-dev set and 83.5% on the Spider-dev set, outperforming other methods.
@@
  • We propose Text2SQL-Flow, a robust Text-to-SQL data augmentation framework that uses existing Text-to-SQL pairs as seeds to systematically generate semantically valid and structurally diverse augmented samples along six dimensions efficiently. An end-to-end automated pipeline is designed, integrating modules for SQL execution filter, NL question generation, prompt generation, CoT generation, and data classification, ensuring high-quality generated data. Furthermore, by introducing a Database Manager module, the framework supports flexible adaptation to various database systems, improving scalability.
  • Based on Text2SQL-Flow, we generate a high-quality Text-to-SQL dataset named SQLFlow, comprising 89,544 samples. The data is efficiently generated, exhibiting high diversity and quality. Experiments on open-source LLMs show that fine-tuning with SQLFlow enhances model performance, achieving competitive results efficiently on benchmark evaluations. Notably, it attained 59.2% on the BIRD-dev dataset and 83.5% on the Spider-test dataset, outperforming other models trained with datasets of similar scale.
@@ II-A Text-to-SQL Techniques
Early rule-based approaches rely on handcrafted templates to map NL questions to SQL queries [li2014constructing, mahmud2015rule, zelle1996learning, an2025unictokens]. Although effective in domain-specific settings, these methods exhibit limited generalizability due to their dependence on rigid rule sets and poor adaptability across diverse database schemas.
The advent of neural sequence-to-sequence architectures marks a significant shift in the field [hochreiter1997long]. Methods such as those in [zhong2017seq2sql, guo2019towards, wang2019rat, chen2021shadowgnn, an2024mc] employ encoder–decoder frameworks augmented with attention mechanisms and schema-aware graph representations to better capture structural relationships between NL and database elements. While these innovations have led to substantial performance improvements, they demonstrate limited robustness when confronted with unseen or heterogeneous schema structures.
The rise of pre-trained language models, such as BERT [devlin2018bert, yin2020tabert] and T5 [raffel2020exploring], further advanced Text-to-SQL research. By combining large-scale unsupervised pretraining with task-specific fine-tuning, these models effectively learn semantic alignments between NL utterances and database schemas [fu2023miga, gu2023few, luo2024llm]. Despite their strong empirical performance, they remain heavily reliant on high-quality, human-annotated training data and are prone to overfitting, especially in low-resource or cross-domain scenarios.
Recently, the emergence of LLMs has opened new directions for Text-to-SQL. Through sophisticated prompting strategies, LLMs can generate accurate SQL queries even in zero-shot or few-shot settings [dong2023c3, liu2023comprehensive]. Techniques including prompt engineering [gao2023text], task decomposition [pourreza2024din], and self-correction have further enhanced their reliability and generalization capabilities, demonstrating considerable promise for cross-domain and real-world applications.
@@ II-B Data-Centric AI for Text-to-SQL
Data-Centric AI for Text-to-SQL can be divided into two main aspects. First is how to better leverage existing data. For example, schema linking structures database schema information to make prompts clearer and reduce ambiguity [nahid2025rethinking, cao2024rsl, li2023resdsql, li2024multisql, lin2025perceiveanythingrecognizeexplain], aiming at improving the model’s understanding of database structures. Additionally, data preprocessing and cleaning [gao2024xiyan, pourreza2025reasoning, glass2025extractive, kothyari2023crush4sql] are crucial, such as constructing inter-table relationship graphs, standardizing field names, and eliminating redundant or conflicting schema information, thereby providing the model with cleaner and more consistent inputs.
[…]
@@ III Methodology
In our work, we propose a data augmentation framework called Text2SQL-Flow, which consists of a data augmentation pipeline and an adapted Database Manager module. Part III-A introduces the data augmentation pipeline, which generates augmented SQL samples from raw seed SQL queries and further synthesizes multidimensional data pairs, as shown in the upper part of Figure 2. Part III-B presents the Database Manager module, as shown in the lower part of Figure 2. Through this framework, we constructed a Text-to-SQL dataset called SQLFlow. The following two parts demonstrate the utilization of SQLFlow. In Part III-C, we use CoT rationales as training targets to supervisedly fine-tune open-source LLMs, thereby enhancing their problem-solving capabilities, as shown in the left part of Figure 3. Part III-D addresses closed-source models and adopts a prompt-based approach. Specifically, we train an embedding model on SQLFlow that aligns the representations of masked NL and SQL queries, thereby improving the effectiveness of the retrieval system, as illustrated in the right part of Figure 3.
III-A Data Augmentation Pipeline
III-A1 SQL Augmentation
Data serves as the cornerstone of Data-Centric AI and plays a pivotal role in training LLMs. However, most accessible data originates from open-source datasets that are limited in scale and stylistically homogeneous. Although proprietary datasets exist, their volume rarely suffices for robust model training. Consequently, data scarcity and uniformity constrain the generalization capacity of LLMs, especially for complex SQL generation tasks.

### v3

@@ Abstract
The data-centric paradigm has emerged as a pivotal direction in artificial intelligence (AI), relying on high-quality training data. This shift is especially critical in the Text-to-SQL task, where the scarcity, limited diversity, and structural simplicity of existing datasets constrain model performance. To address these challenges, we propose Text2SQL-Flow, a SQL-aware data augmentation framework that systematically generates large-scale, semantically valid, and structurally diverse Text-to-SQL pairs from limited seed data. Our framework operates along six augmentation dimensions and integrates an end-to-end pipeline featuring auxiliary database selection, SQL executability verification, natural language (NL) question generation, NL–SQL correspondence verification, and chain-of-thought (CoT) reasoning trace generation. Leveraging this framework, we construct SQLFlow, a high-quality dataset comprising 75,386 annotated examples. We demonstrate the utility of SQLFlow in both fine-tuning and prompt-based settings: (1) For open-source large language models (LLMs), fine-tuning with SQLFlow enhances the problem-solving capabilities. Under the same data budget, models trained on SQLFlow achieve competitive performance gains across multiple benchmarks. (2) For closed-source LLMs, we propose a masked alignment retrieval method that leverages SQLFlow as both a knowledge base and the training data for the retrieval model. This approach enables structure-aware example matching by modeling fine-grained alignments between NL questions and SQL queries. Experimental results show that our retrieval strategy outperforms existing example retrieval methods, highlighting the dual importance of SQLFlow’s high-quality data and our novel retrieval technique. Our work establishes a scalable, data-centric foundation for advancing Text-to-SQL systems and underscores the indispensable role of structured, high-fidelity data in modern AI development. Our code is available at https://github.com/TechNomad-ds/Text2SQL-Flow.
@@ I Introduction
In recent years, the data-centric artificial intelligence (AI) paradigm has garnered increasing attention [54, 22]. Unlike traditional algorithm-centric approaches that primarily focus on advancing model architectures and algorithms, many AI applications are increasingly constrained by the availability of high-quality data. Vast amounts of data remain underutilized, despite their immense potential value. From a data-centric perspective, high-quality, diverse, and well-structured training data plays a central role in advancing AI [18].
@@
Window Func. Set Op. Subquery GROUP BY # CASE # WHERE # JOIN
Spider-train 8,659 0.00 6.07 15.59 22.60 0.00 0.86 0.71
SQLFlow-Spider 30,502 14.37 9.49 40.06 37.67 0.22 1.84 0.96
@@
SQLFlow-EHRSQL 11,721 21.65 7.47 66.86 34.14 0.16 4.32 0.95
SQLFlow 75,386 14.63 8.11 42.57 33.68 0.23 2.46 1.02
@@
The Text-to-SQL task is currently facing an urgent demand for high-quality data. Although the structured query language (SQL) enables powerful and precise database interactions, its complex syntax and steep learning curve limit its accessibility for non-expert users. The Text-to-SQL task aims to translate natural language (NL) questions into executable structured query language (SQL) queries, allowing users without specialized knowledge to access databases conveniently [23]. This capability shows broad application prospects in intelligent analysis, business intelligence platforms, and enterprise data services. With the rise of LLMs, significant progress has been made in Text-to-SQL research.
Figure 1: Performance comparison of models trained on SQLFlow vs. SynSQL [29] (both at 75,386 samples). Blue labels indicate results for fine-tuned open-source models, while green labels indicate results of closed-source model with few-shot retrieval for prompt construction.
@@
Existing methods are mainly divided into two categories: (1) fine-tuning-based methods [10, 11, 31, 41] and (2) prompt-based methods [45, 42, 30, 38]. In both categories of methods, high-quality data plays a pivotal role. Fine-tuning-based methods rely on data pairs to enable LLMs to learn the mapping from NL questions to SQL queries, thereby enhancing the model’s reasoning capability. In prompt-based methods, where model weights are inaccessible, in-context learning becomes key to improving performance, and the quality of retrieved few-shot examples is critical. Data can serve as reference examples in a knowledge base or be used to train retrieval models for obtaining more representative few-shot samples [33, 37, 10].
Since the Text-to-SQL task relies on NL questions and corresponding SQL queries, efficiently acquiring large-scale, high-quality data remains a fundamental challenge. Manual collection not only struggles to guarantee quality but is also limited in scale. Existing public datasets (e.g., Spider [52], BIRD [32]) mainly rely on manual annotation, which leads to limited size and narrow data coverage. Recent studies [47, 29] have explored the potential of automatically synthesizing Text-to-SQL datasets at scale using LLMs. However, these approaches often overlook the rich structural and semantic information already embedded in existing curated Text-to-SQL datasets. These datasets, which are typically carefully curated and execution-verified, provide a strong foundation for systematic and scalable data expansion. Failing to fully leverage such high-quality resources may result in suboptimal data efficiency and unnecessary redundancy in data generation.
Data diversity in Text-to-SQL can be characterized along two dimensions. Distribution-level diversity refers to variations in underlying database schemas, data domains, and result distributions, and is typically introduced by incorporating databases from different sources or domains. In contrast, query-level diversity captures the structural and compositional variations of SQL queries, such as aggregation patterns, join complexity, nested subqueries, and logical constraints.
To address these challenges, we propose Text2SQL-Flow, a robust and scalable Text-to-SQL data augmentation framework that systematically expands existing seed data derived from original databases and their corresponding SQL queries. To enhance distribution-level diversity, our framework first introduces auxiliary databases by selecting databases with varying structural complexities and data domains, thereby enriching the overall data landscape. Building upon seed SQL queries from both original and auxiliary databases, Text2SQL-Flow defines six data augmentation dimensions spanning data value transformations, query structure modifications, and complexity enhancements. Leveraging LLMs, the framework synthesizes diverse and semantically valid augmented SQL queries. A series of specialized operators is then employed to perform SQL executability filtering, NL question generation, NL–SQL alignment filtering, prompt construction, and chain-of-thought (CoT) reasoning path generation. These operators collaboratively ensure the quality, correctness, and diversity of the generated data. Through the combination of LLM-driven generation and automated filtering mechanisms, Text2SQL-Flow is capable of producing large-scale, high-quality, and diverse Text-to-SQL datasets from limited original data, while maintaining strong scalability and extensibility across heterogeneous databases. Furthermore, a unified database interaction layer is adopted to abstract database-specific details, enabling efficient SQL execution and schema access across multiple relational database systems.
@@
Based on our SQL augmentation pipeline, a high-quality Text-to-SQL dataset comprising 75,386 entries called SQLFlow is constructed. Using SQLFlow, we explore the applications of data in Text-to-SQL tasks. For open-source LLMs, the augmentation framework can provide higher-quality training data. The included CoT can further assist models in improving their problem-solving capabilities. After fine-tuning open-source models with data generated by our framework, we achieved excellent performance on multiple benchmarks, validating the framework’s effectiveness. As shown in Figure 1, under the same data scale, the model trained by SQLFlow achieves higher performance than SynSQL [29] across all benchmarks, attaining 56.8% on the BIRD-dev dataset. For closed-source models, we adopt prompt-based methods, retrieving few-shot examples to provide references for the LLM in solving problems, thereby achieving broad applicability across various scenarios. Existing systems typically select examples from limited public datasets, which have significant limitations in semantic diversity and structural coverage. Therefore, the knowledge base for retrieval can also be constructed using the data generated by our framework. However, at the retrieval method level, current mainstream retrieval strategies often focus merely on surface-level semantic similarity, neglecting the deep syntactic and logical relationships between NL questions and SQL queries. To address this, we propose an improved masked alignment retrieval method for example selection. This method learns the masked alignment relationships between NL questions and SQL queries during the training phase and achieves fine-grained structure-aware matching during the inference phase, enhancing the relevance and effectiveness of retrieved examples. The training data for this retrieval model also utilizes the data constructed by our framework. As shown in Figure 1 shows that our retrieval method achieves competitive performance on benchmarks, with 61.0% on the BIRD-dev and 83.5% on the Spider-dev, outperforming other methods.
@@
  • We propose Text2SQL-Flow, a robust Text-to-SQL data augmentation framework that uses existing Text-to-SQL pairs as seeds to systematically generate semantically valid and structurally diverse augmented samples along six dimensions efficiently. An end-to-end automated pipeline is designed, integrating modules for auxiliary database selection, SQL augmentation generation, SQL executability filter, NL question generation, NL-SQL alignmentation filter, prompt generation, and CoT generation, ensuring high-quality generated data.
  • Based on Text2SQL-Flow, we generate a high-quality Text-to-SQL dataset named SQLFlow, comprising 75,386 samples. The data is efficiently generated, exhibiting high diversity and quality. Experiments on open-source LLMs show that fine-tuning with SQLFlow enhances model performance, achieving competitive results efficiently on benchmark evaluations. Notably, it attained 61.0% on the BIRD-dev dataset and 85.1% on the Spider-test dataset, outperforming other models trained with datasets of similar scale.
@@ II-A Text-to-SQL Techniques
Early rule-based approaches map NL questions to SQL using handcrafted templates [28, 34, 53], but their reliance on rigid rules severely limits generalization across domains and schemas. Neural sequence-to-sequence models [57, 17, 43, 3] introduce encoder–decoder architectures with attention and schema-aware representations to model the alignment between NL and database structures. However, these methods still struggle with robustness when generalizing to unseen or heterogeneous schemas. Pre-trained language models such as BERT [5, 49] and T5 [40] further improve Text-to-SQL by enhancing semantic alignment between NL utterances and database schemas [7, 14]. Nonetheless, their performance heavily depends on large amounts of annotated data and degrades in low-resource or cross-domain settings.
Recently, the emergence of LLMs has opened new directions for Text-to-SQL. Through sophisticated prompting strategies, LLMs can generate accurate SQL queries even in zero-shot or few-shot settings [6, 33]. Techniques including prompt engineering [10], task decomposition [38], and self-correction have further enhanced their reliability and generalization capabilities, demonstrating considerable promise for cross-domain and real-world applications.
@@ II-B Data-Centric AI for Text-to-SQL
Data-centric approaches for Text-to-SQL focus on leveraging existing data and synthetic data generation. For existing data, prior work improves schema understanding via schema linking and structural representations [36, 2, 30, 27], along with preprocessing and cleaning techniques [11, 39, 12, 25] that refine schema information and inter-table relationships.
Synthetic data generation methods are generally divided into SQL-to-question and question-to-SQL paradigms. SQL-to-question approaches [15, 44, 46, 56] generate natural language questions from SQL queries using templates or rules, producing high-fidelity pairs but suffering from limited scalability and difficulty in handling complex queries. Question-to-SQL methods [48, 45, 50, 51] generate questions first and then infer SQL using Text-to-SQL models; while more natural, these methods are prone to error propagation. Recent work seeks to mitigate these issues. Sense [47] adopts prompt chaining to jointly generate schema, question, and SQL with semantic consistency, but relies on costly closed-source models. OmniSQL [29] instead decouples the generation pipeline and leverages open-source or lightweight LLMs with CoT reasoning, offering a more scalable and reproducible solution.
@@ III Methodology
Despite the growing interest in the Text-to-SQL task, its development is fundamentally constrained by the limited availability of high-quality SQL data. For a given Text-to-SQL system, the user typically provides an existing database $d_{o}$, which serves as both the available database resource and the target database on which the Text-to-SQL task is performed. Although there exists a set of manually annotated SQL queries $\mathcal{Q}_{o}$ over $d_{o}$, its scale is usually limited. The objective is therefore to improve the model’s performance on the Text-to-SQL task over $d_{o}$. However, due to the high cost of expert annotation, $\mathcal{Q}_{o}$ is small in scale, making it insufficient to rely solely on it for learning robust and generalizable query patterns. Nevertheless, $\mathcal{Q}_{o}$ is typically of high quality, with accurate semantics and strong alignment to the underlying database schema, making it a reliable foundation for data expansion. As a result, how to effectively leverage $\mathcal{Q}_{o}$ to construct large-scale and diverse SQL data remains challenging.
When expanding the scale of SQL data, diversity should not be overlooked. SQL diversity can be characterized along two complementary dimensions. Distribution-level diversity captures variations in database schemas, data domains, and underlying data distributions. Such diversity enhances domain coverage and mitigates overfitting to database-specific patterns. Meanwhile, query-level diversity reflects structural and compositional variations of SQL queries. Both dimensions are crucial for constructing high-quality augmented SQL datasets. Motivated by these insights, we propose a joint data augmentation framework, Text2SQL-Flow (shown in Figure 2), which simultaneously enhances distribution-level diversity and query-level diversity under data-scarce settings. Based on this framework, we further design a data generation pipeline to construct large-scale, high-quality, and diverse SQL data.
III-A Distribution-level Diversity Augmentation
Distribution-level diversity captures variations in database schemas, data domains, and underlying data distributions. When the original database $d_{o}$ is fixed, an effective way to improve distribution-level diversity is to introduce additional databases as auxiliary resources. Each auxiliary database is typically accompanied by its own set of existing SQL queries, which naturally expands the diversity of SQL patterns. Moreover, newly introduced databases bring novel data domains and distinct data distributions, thereby enriching the space of feasible SQL augmentations. This is because SQL queries are inherently shaped by the underlying data distributions they operate on. However, the number of available databases is often large, making it impractical to incorporate all of them. We therefore construct a database warehouse $\mathcal{D}_{w}$ consisting of publicly available and accessible databases. Each database $d\in\mathcal{D}_{w}$ is associated with a corresponding SQL set $\mathcal{Q}(d)$. Given resource constraints, it is infeasible to utilize the entire warehouse. Instead, we select a subset of databases from $\mathcal{D}_{w}$ as auxiliary databases, denoted as $\mathcal{D}_{a}$, to enhance distribution-level diversity. Our key principle for selecting $\mathcal{D}_{a}$ is that, under a limited budget, the chosen auxiliary databases should cover a wide range of schema complexities.
For each candidate database $d\in\mathcal{D}_{w}$, we compute a set of complexity metrics that depend solely on the database schema structure, including the number of tables $T(d)$, the total number of columns $C(d)$, and the foreign key density $\text{FK\_density}(d)=\frac{\text{FK}(d)}{\max(T(d),\epsilon)}$ where $\epsilon$ is a small smoothing constant used to ensure numerical stability when the number of tables is small. To avoid inconsistencies in scale across different metrics and the need for manually specified weights, we compute the percentile rank of each complexity metric over the database warehouse $\mathcal{D}_{w}$. Specifically, for any metric $x(d)$, its percentile rank is defined as $r_{x}(d)=\frac{1}{|\mathcal{D}_{w}|}\sum_{d^{\prime}\in\mathcal{D}_{w}}\mathbb{I}\!\left[x(d^{\prime})\leq x(d)\right]$ where $\mathbb{I}[\cdot]$ denotes the indicator function. Accordingly, we obtain $r_{T}(d)$, $r_{C}(d)$, and $r_{FK}(d)$. Based on these rankings, we define the schema complexity score of a database as
$\text{score}(d)=r_{T}(d)+r_{C}(d)+r_{FK}(d).$ (1)
We then sort all candidate databases according to $\text{score}(d)$ and select $k$ quantile points at uniform intervals over the score distribution, thereby determining $k$ auxiliary databases.
We then construct a combined corpus consisting of the SQL queries from the original database $\mathcal{Q}_{o}$ and those from the auxiliary databases $\bigcup_{d\in\mathcal{D}_{a}}\mathcal{Q}(d)$. Although this process substantially increases distribution-level diversity, the SQL queries themselves remain constrained by the structural patterns inherent in the original datasets. Therefore, an augmentation strategy over existing SQL queries is required to further expand query-level diversity.
III-B Query-level Diversity Augmentation
III-B1 SQL Augmentation